Petrol​ Ibérico. Petrol​ Ibérico, a European gas​ company, is borrowing ​$750,000,000 via a syndicated eurocredit for six years at 70 basis points over LIBOR. LIBOR for the loan will be reset every six months. The funds will be provided by a syndicate of eight leading investment​ bankers, which will charge​ up-front fees totaling 1.3​% of the principal amount. What is the effective interest cost for the first year if the annual LIBOR is 3.50​% during the first six months and 3.90​% during the second six months.

The effective interest cost for the first year is %. ​(Round to two decimal​ places.)
Answer #1

Net proceeds are borrowed amount less of up-front fees of 1.3%

Net proceeds = 750,000,000 - 1.3% of 750,000,000 = 740,250,000

Interest for the first 6 months

= Net proceeds * LIBOR + Above Libor charge / 2

=740,250,000 * (3.50% + 0.7%)/2

=15,545,250

Interest for the second 6 months

= Net proceeds * LIBOR + Above Libor charge / 2

=740,250,000 * (3.90% + 0.7%)/2

=17,025,750

Total interest = 15,545,250 + 17,025,750 = 32,571,000

Effective interest rate

= Total interest / Net Proceeds

=32,571,000 / 740,250,000

=0.044

It is for 6 months, the annual effective rate would be 0.044*2

=8.8%